The population P (in millions) of California x years after 2000 can be modeled by P=34e^0.013x.
a.) Use properties of logarithms to solve this equation for x.
b.) Use your equation to find x when P=38.Interpret your answer.

Answer by nerdybill(7384)

The population P (in millions) of California x years after 2000 can be modeled by P=34e^0.013x.
a.) Use properties of logarithms to solve this equation for x.
P=34e^(0.013x)
P/34 = e^(0.013x)
ln(P/34) = 0.013x
ln(P/34)/0.013 = x
.
b.) Use your equation to find x when P=38.Interpret your answer.
ln(P/34)/0.013 = x
ln(38/34)/0.013 = x
.11123/0.013 = x
8.56 = x
.
Year: 2000+8.56 = 2008.56
The population of California will be 38 million around mid-year of 2008.